OpenStudy (anonymous):

Will Give Medal!!! Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3 A. y - 7 = - 2/3 (x + 1) B. y + 7 = 2/3 (x + 1) C. y + 7 = - 2/3 (x - 1) D. y - 7 = 2/3 (x - 1)

OpenStudy (anonymous):

OMG i love doing slope do you know the slope equation

OpenStudy (anonymous):

I can't figure out which one is it. I'm not too good at Slope..

OpenStudy (anonymous):

ok so M=y^2-y^1 divded by x^2 - x^1

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

so the point 1 aand -7 would equal the y intercept so y-7 right

OpenStudy (imtiaz7):

here we use point slope form

OpenStudy (imtiaz7):

that is y-y1=mx-x1 where x1=1 , y=-7 and m=-2/3

OpenStudy (imtiaz7):

y-y1=m(x-x1) so by putting the valuse we get the equation: so c is the right answer :)

OpenStudy (anonymous):

Thank You! Thanks for explaining it all too.

OpenStudy (imtiaz7):

@sarahstyles welcome..

OpenStudy (anonymous):

I gave you a medal (as promised) and a follow :)

OpenStudy (imtiaz7):

... if you'll not do that then i'll still help you..

OpenStudy (anonymous):

Thanks :)

OpenStudy (imtiaz7):

:)